Suppose Maria gets R13240.10 after investing a lump for 5 years.The bank quoted 11.5% simple interest.How much did Maria invest at the beginning?


1
Expert's answer
2021-06-07T14:33:46-0400



   Findings from the question:

Amount received after five years = R 13,240.10 (future value )



Simple interest rate "= 11.5\% = \frac{11.5}{100 }= 0.115"

Time = 5 years 

Amount invested in the beginning = ? (present value)


"Present \space Value=\frac{Future \space Value}{1+rt}"


"=\frac{\$13,240.10}{(1+0.115\times)}"


"=\frac{\$13,240.10}{1.575}"


"=\$8,406.41"


At the beginning, Maria had invested $ 8,406.41 on which she gets lump sum amount of $13,240.10.